OpenStudy (anonymous):

EXPRESS IN SIMPLEST RADICAL FORM. ALL VARIABLE REPRESENT POSITIVE REAL NUMBERS

OpenStudy (anonymous):

\[4/5 \sqrt{72XY ^{3}}\]

OpenStudy (anonymous):

You can break this up and take the square roots seperately. It can be expressed as \[4/5 * \sqrt{72} *\sqrt{x} *\sqrt{y ^{3}}\] \[4/5*6*\sqrt{2}*\sqrt{x} * y \sqrt{y}\] this can be simplified to \[24y/5* \sqrt{2xy}\]